OpenStudy (anonymous):

what are the different combinations of 6C3

OpenStudy (mathmale):

If all you want is to evaluate \[_{6}C _{3}\] that is relatively easy (assuming that you know and understand the "factorial" operator ! ). Note that 3! = 3 * 2 * 1, 4! = 4 * 3 * 2 * 1, and so on. \[_{6}C _{3}=\frac{ 6! }{ 3!(6-3)! }\]

OpenStudy (mathmale):

Give that a try, OK?

OpenStudy (anonymous):

thank you so much:)

OpenStudy (mathmale):

You're very welcome, Hannah!! Good luck!